# Benchmark the Linguistic Competence of Language Models

### v1

@@ 3.3 Summary
This meta-study emphasizes the need to consolidate existing resources for a comprehensive assessment of the linguistic competence of LMs — a manifold but rather blind spot in evaluation research. Apart from more thorough evaluations, such a stimulus can significantly boost future research, as happened in computer vision with ImageNet (Deng et al., 2009) or in NLP with GLUE and SuperGLUE (Wang et al., 2019a, b).
@@ 4.1 Datasets
To feature a total of 208 unique datasets, we leverage existing and established resources like OntoNotes (Weischedel et al., 2013), English Web Treebank (Silveira et al., 2014), or BLIMP (Warstadt et al., 2020) and create datasets addressing phenomena like the POS of words, their dependencies or determine the linguistic acceptability of sentences. Further, we include a range of less employed data, addressing contextualization of words (Klafka and Ettinger, 2020), reasoning (Talmor et al., 2020), semantic decomposition (White et al., 2016; Rudinger et al., 2018a, b; Govindarajan et al., 2019; Vashishtha et al., 2019), grammatical knowledge (Huebner et al., 2021), bridging (Pandit and Hou, 2021), and rhetorical (Carlson et al., 2001) and discourse (Webber et al., 2019) structure in text. Finally, we cover rarely probed phenomena like negation (Szarvas et al., 2008; Konstantinova et al., 2012; Vahtola et al., 2022), or word complexity (Paetzold and Specia, 2016).
@@ 4.2 Structure
Apart from the comprehensive scope, Holmes provides a clear structure for specific evaluations on different levels of aggregation. We first group the datasets according to the linguistic phenomena addressed. Then, we categorize these phenomena into their previously introduced type (see § 2) - morphology, syntax, semantics, reasoning and discourse. We rely on the categorization provided by the specific studies whenever given. The detailed categorization is given in § A.3.
@@ 4.3 Experimental Setup
Holmes evaluation follows the primarily used classifier-based probing paradigm, as described in § 2. Considering the internal representations allows us to maximally disentangle the understanding of distinct linguistic phenomena from each other and from other cognitive abilities (like following textual instructions). Further, this method allows us to assess any type of LMs, including sparse, static, or contextualized ones. Based on the specific dataset, we either select the embeddings of the specific input tokens (like single words for POS tagging) or average embeddings across a span or the whole sentence. We define a probing task as training a probe $f_{p}$ (linear model without intermediate layers) using these embeddings as inputs and the dataset labels as training signals. If not defined in the original data, we divide the dataset samples into train/dev/test split following a ratio of 70/10/20. We repeat this procedure five times using different random seeds and aggregate the results afterward.
@@ 4.4 Evaluations
Next, we evaluate the probes’ reliability from an information-theoretic perspective. Following Voita and Titov (2020), we use the compression $I$ to measure how well a probe compresses input data. A higher $I$ means fewer bits are needed, indicating that the given linguistic phenomenon is more clearly encoded in the embeddings.
@@
A reliable probe should grasp patterns relevant to the tested phenomena in the internal representations of LMs but should not be able to learn anything else. Therefore, we expect high performance when evaluating the specific dataset but low performance when we randomize training signals. We check this using control tasks introduced in Hewitt and Liang (2019). Specifically, we calculate the selectivity $S$ as the difference between the probe trained with the original labels $y$ and the control task where we train the probe with randomly assigned labels $y^{\prime}$. With a higher $S$, we assume the detected patterns are relevant for the specific phenomena under test, as random patterns do not lead to similar performance.
@@ 5 Holmes Results
Using Holmes, we evaluate a diverse collection of 59 LMs. [^2] Find a complete list in Appendix § A.2. Using the results of these extensive experiments, we first answer the research question: what is the linguistic competence of LMs? In doing so, we discuss the reliability of results (i), the linguistic competence of LMs concerning the unique structure of Holmes (ii), and how these results relate to other downstream abilities (iii). Subsequently, we examine how linguistic competence varies among LMs, as we find LMs prevailing for different types of linguistic phenomena (Figure 1) and delve into the effects of model architecture (iv), size (v), and instruction tuning (vi).
i) The reliability of Holmes.
First, we show the reliability of probing-based evaluation, using deviation $\sigma$, compression $I$, and selectivity $S$ results in Figure 7. Single outliers are datasets that are too hard for all LMs, as the sample size is too small, or the linguistic phenomena under test are too complex, as the ability to detect spans causes speculations in a text. We average these metrics for every dataset across all LMs. Note, for selectivity, we consider only base-sized model (10m-200m parameters) for computational efficiency.
@@
First, we found a low average deviation ($\sigma=0.02$), indicating the high reliability of probes across random seeds. These results also highlight the stability of probing results, compared to prompting-based ones where results across many paraphrased prompts lead to a deviation of $\sigma=0.07$ reported in Mizrahi et al. (2024). Next, substantial compression (average $I=1.9$) and selectivity (average $S=0.31$) further confirm the probes’ reliability. Interestingly, one identifies two parallel trends for selectivity. Harder datasets with many labels, like POS tagging, are arranged around a selectivity of 0.1 to 0.4 and a task metric of 0.3. In contrast, for easier binary classification tasks (such as linguistic applicability), we observe selectivity around 0.2 to 0.5 and a task metric of 0.6 to 0.9. Further, we measure a significant ($p<0.05$) positive correlation between the task metrics and the compression ($\tau=0.64$) and selectivity ($\tau=0.65$). This further confirms our reliability assumption and allows us to trust the task metric as the primary evaluation measure.
@@
Figure 7: Reliability evaluation using deviation, compression (log), and selectivity on the y-axis for all 208 probing datasets. The x-axis represents the task metrics (either person correlation or macro $F_{1}$).
ii) The story of Holmes.
@@
We focus on what Holmes tells us in general and regarding formal and functional phenomena, as defined in § 2. We report in Figure 8 the task metric, discriminability, and selectivity, averaged for every phenomena type. Note, discriminability (Rodriguez et al., 2021) quantifies the alignment of LMs ranking of one specific dataset compared to the overall rankings using the Kendall Tau correlation. Considering these three metrics, all tested LMs strongly encode formal phenomena (morphology and syntax), which often depend on the local neighborhood of words. Therefore, we assume that LMs approximate these co-occurrences during pre-training with high precision. For example, the specific POS tag of a word, like man (noun), primarily depends on its surroundings, such as the frequent predecessor the. In contrast, LMs encode less information about functional phenomena (semantics, reasoning, and discourse) since they show a relatively low performance regarding the task metric. For these functional phenomena, we assume more complex co-occurrences are required to capture the broad context in language, such as the rhetorical relation of two distant text spans. Despite these differences between formal and functional phenomena types, they contribute to the benchmark in a balanced way. A low to medium discriminability indicates that none of these types of linguistic phenomena dominates the overall LM rankings.
@@
This balanced influence of the five phenomena types is further visible when considering their ranking correlations (Figure 9, left). A high average correlation of $67.8\pm 6.6$ with the overall results (last column/row) hints that they are facets of a broader occurrence but share common characteristics. Still, breaking into categories is meaningful, as the phenomena types (first five columns/rows) are medium correlated (average of $53.9\pm 14.5$). Analyzing the results of phenomena types further highlights the value of this distinction. While results of morphology and syntax are similarly correlated with the overall results ($68.2$ and $70.2$), their direct correlation ($69.1$) indicates their supplementary nature. Further, discourse results show the lowest correlation with others ($44.8\pm 16.1$), indicating the particular scope.
@@
iii) The companions of Holmes.
We analyze how the results of Holmes and those from other evaluations focusing on downstream applications align (Figure 9, right). We select the OpenLLM benchmark (Beeching et al., 2023), as it covers a wide range of open LMs, in contrast to others like HELM (Liang et al., 2023). First, Holmes and OpenLLM results of jointly evaluated LMs are medium correlated, hinting that the linguistic competence of LMs is partly aligned with their downstream abilities. The nature of this alignment is further evident when focusing on morphology, reasoning, and discourse. Interestingly, and in contrast to syntax and semantics, their correlation to the OpenLLM and Holmes overall results is similar. Therefore, these three phenomena presumably represent skills that are more tested in the general benchmarks. These correlation patterns are consistent across the three most meaningful OpenLLM datasets (MMLU, TruthfulQA, and GSM8K). As TruthfulQA shows lower correlations with the linguistic phenomena and other datasets within OpenLLM, we presume this dataset captures distinctly different skills (possibly knowledge). [^3] Further, it’s also known that we need to expect this dataset to be fully leaked (Balloccu et al., 2024). These insights show how different benchmarks provide a different scope and supplement themselves simultaneously. Further, the above analysis shows, again, the value of assessing the linguistic competence of LMs across different phenomena types, for fine-grained analyses.
Figure 9: Kendall-tau correlation within Holmes (left) and compared to the OpenLLM benchmark (right). Green stars indicate significant correlations ($p<0.05$).
@@
iv) The effect of language model architecture.
Next, we discuss the impact of model architecture on the linguistic competence of LMs. In Figure 11 (left), we compare encoder and decoder LMs. Due to the absence of big encoder LMs, we consider five encoder and six decoder LMs with up to 220m parameters. Encoder LMs show a higher $mwr$ of $52\%$ than decoder LMs ($21\%$). This observation is the most saturated for morphology or syntax, encompassing a variety of token-level phenomena, like part-of-speech. We assume that the missing bi-directional encoding of decoder LMs causes this lower performance because the available context of one token heavily depends on its position. Thus, even common tokens, like the, have different potential representations - at the beginning or in the middle of a sentence. These instabilities are further evident when considering Figure 11 (right) which reports the accuracy for the top-20 most common POS tokens (such as the) based on the pos, xpos, upos dataset. Given their high frequency, one expects stable prediction performance. Surprisingly, encoder LMs (BERT and RoBERTa) show higher median accuracy and clearly lower deviations compared to the same-size decoder counterpart (GPT2). While scaling model size to 12B (Pythia) and 70B (Llama-2) allows for improved accuracy and lower deviations, decoder LMs do not match the encoder performance, even up to 700 times bigger.
v) The effect of scaling parameters.
We discuss how the number of parameters influences the linguistic competence of LMs. Given the variety of LMs of different sizes, we focus on the Pythia (decoder-only) and T5 (encoder-decoder) families. From Figure 10, we observe for both Pythia and T5 that the linguistic competence scales with model size, and it is particularly pronounced after exceeding 0.5B (Pythia) and 1.0B (T5) parameters. Again, model architecture is crucial, as T5 LMs (encoder-decoder) exhibit a clearly higher mean winning rate of $40-70\%$ than Pythia (decoder-only) ones with $mwr$ of $20-60\%$. Further, we found formal phenomena evolving differently with increased model size than functional ones. Specifically, morphology and syntax start at a lower level, with an apparent performance jump after 0.5B (Pythia) and 1.0B (T5) parameters, followed by slow but steady growth. Differently, semantics, reasoning, and discourse start at a higher $mwr$, followed by a continuous improvement as the model size grows. From these results, we assume more parameters allow LMs to better approximate simpler co-occurrences in the near neighborhood of words to understand formal phenomena like word dependencies. In contrast, more parameters do not have the same pronounced effect on functional phenomena, like rhetorical relations, which require an LM to acquire more distant and complex word co-occurrences.
@@
Tülu-2 +5% +2% -15% 0% -30% -8%
Orca-2 -1% -3% -4% +4% -5% -2%
Llama-2-chat +3% +1% -6% +3% -1% -1%
Vicuna-v1.5 +23% +7% -3% +6% -6% +4%
@@
Tülu-2 +15% 0% -11% -3% 0% -2%
Llama-2-Chat +23% +14% +2% +4% +17% +10%
Average +10% +4% -3% +4% -2% +1%
@@
Table 1: Effect of instruction tuning on the mean winning rate compared to the pre-trained LMs.
vi) The effect of instruction tuning.
@@
Finally, we focus on how instruction tuning affects LMs’ linguistic competence and compare the tuned LMs with their base models—for example, FLAN-UL2 vs. UL2. From results in Table 1, we note less saturated effects for the overall scope while being more pronounced for the five phenomenon types - again emphasizing the structured and comprehensive evaluation of linguistic competence. On average, we found instruction tuning has the highest effect on morphology ($+10\%$) followed by syntax ($+4\%$), reasoning ($+4\%$), and a negative effect for semantics $-3\%$ and discourse $-2\%$. These results confirm previous assumptions that instruction tuning updates are often superficial (Yadav et al., 2023; Hershcovitch et al., 2024; Sharma et al., 2023) and that LMs are better at mimicking language (formal phenomena) than understanding it, measured with functional phenomena (Mahowald et al., 2024). Further, larger models benefit more from instruction tuning. Llama-2-70b-Chat and FLAN-UL2 gain up to $+23\%$ and $+40\%$ for morphology and $+10\%$ and $+13\%$ on average. In addition, decoder-only LMs (Llama-2 and Pythia) tend to show less pronounced positive effects than encoder-decoder LMs (FLAN-T5-XXL and FLAN-UL2). However, they better understand reasoning phenomena. When comparing LMs based on Llama-2-13b, we see that specific fine-tuning methods shape the LMs differently. The top-ranked 13b LM for Holmes and OpenLLM, Vicuna, was trained on 125k instructions, less than other models. Thus, high quality is more important than the number of instructions for LMs’ linguistic competence. Tülu loses performance while being trained on a large mixture of data (approx. 330k instructions), the same for its 70b version. Finally, the focus of Orca-2 on reasoning is also reflected in its embedding space. These insights show again that while providing a particular perspective, Holmes shows clear differences between LMs and allows us to map them to methodological decisions.

### v4

@@ 3.3 Summary
Our meta-study emphasizes the need to consolidate existing resources for a comprehensive assessment of the linguistic competence of LMs — a manifold but rather a blind spot in evaluation research. Apart from more thorough evaluations, such a stimulus can significantly boost future research, as happened in computer vision with ImageNet (Deng et al., 2009) or in NLP with GLUE and SuperGLUE (Wang et al., 2019a, b).
@@ 4.1 Datasets
We provide a comprehensive coverage of linguistic phenomena by covering 208 unique datasets. We leverage existing and established resources like OntoNotes (Weischedel et al., 2013), English Web Treebank (Silveira et al., 2014), or BLiMP (Warstadt et al., 2020) to create datasets addressing phenomena like the POS of words, their dependencies or the linguistic acceptability of sentences. Further, we include a range of less employed data, addressing contextualization of words (Klafka and Ettinger, 2020), reasoning (Talmor et al., 2020b), semantic decomposition (White et al., 2016; Rudinger et al., 2018a, b; Govindarajan et al., 2019; Vashishtha et al., 2019), grammatical knowledge (Huebner et al., 2021), bridging (Pandit and Hou, 2021), and rhetorical (Carlson et al., 2001) and discourse (Webber et al., 2019) structure in text. Finally, we cover rarely probed phenomena like negation (Szarvas et al., 2008; Konstantinova et al., 2012; Vahtola et al., 2022), or word complexity (Paetzold and Specia, 2016).
@@ 4.2 Structure
Apart from the comprehensive scope, Holmes provides a clear structure for specific evaluations on different levels of aggregation. We first group the datasets according to the linguistic phenomena addressed. Next we categorize these phenomena into their previously defined five phenomena types (see § 2): morphology, like the agreement of subject and verb; syntax, such as the part-of-speech of words; semantics, like semantic roles of words; reasoning, such as detecting a negated sentence; and discourse, like selecting the correct following sentence. Table 1 provides examples for every type of phenomenon. Note that we rely on the categorization provided by the specific studies whenever given - more details in the Appendix § A.3. For example, Conneau et al. (2018) categorized the tense of the main clause as semantic. This phenomenon could also be categorized as syntax if we test the detection of incorrect formulations given a specific tense. However, we follow the authors’ suggestion and test the detection of the tense on a sentence level, which represents semantic aspects.
@@ 4.3 Experimental Setup
Holmes evaluation follows the primarily used classifier-based probing paradigm, as described in § 2, to analyze the internal representations of the last layer of LMs. [^2] Please refer to Appendix § A.4 and Figure 14 for more details about the composition of the internal representations. Thereby, we maximally disentangle the understanding of distinct linguistic phenomena from each other and from other cognitive abilities, such as following textual instructions. Further, this method allows us to assess any LM tpye, including sparse, static, or contextualized ones. Based on the specific dataset, we either select the embeddings of the specific input tokens (like single words for POS tagging) or average embeddings across a span or the whole sentence. We define a probing task as training a probe $f_{p}$ (linear model without intermediate layers) using these embeddings as inputs and the dataset labels as training signals. If not defined in the original data, we divide the dataset samples into train/dev/test split following a ratio of 70/10/20. We repeat this procedure five times using different random seeds for a robust measurement.
@@ 4.4 Evaluations
Next, we evaluate the probes’ reliability from an information-theoretic perspective. Following Voita and Titov (2020), we use the compression $I$. It is the ratio between the minimum description length $mdl$ of encoding $n$ instances with a label space of $K$ compared to applying a uniform encoding $I=\dfrac{u}{mdl}$. A higher $I$ means fewer bits are needed to encode the instance and their labels, indicating that the given linguistic phenomenon is more clearly encoded in the internal representation of LMs.
@@
A reliable probe should grasp patterns relevant to the tested phenomena in the internal representations of LMs but should not be able to learn anything else. Therefore, we expect high performance when evaluating the specific dataset but low performance when we randomize training signals. We check this using control tasks introduced in Hewitt and Liang (2019). Specifically, we calculate the selectivity $S=F_{1}(y,\hat{y})-F_{1}(y^{\prime},\hat{y^{\prime}})$ as the difference between the probe trained with the original labels $y$ and the control task where we train the probe with randomly assigned labels $y^{\prime}$. With a higher $S$, we assume the detected patterns are relevant for the specific phenomena under test, as random patterns do not lead to similar performance.
@@ 5 Holmes Results
Using Holmes, we evaluate a diverse collection of 59 LMs. [^3] Please refer to Appendix § A.2 for a complete list. Using the results of these extensive experiments, we first answer the research question: what is the linguistic competence of LMs? In doing so, we discuss the reliability of results (i) and the linguistic competence of LMs concerning the unique structure of Holmes (ii). Subsequently, we examine how linguistic competence varies among LMs, as we find LMs prevailing for different types of linguistic phenomena (Figure 7) and delve into the effects of model architecture (iii), size (iv), and instruction tuning (v). Finally, we show how Holmes’ results relate to the linguistic performance of LMs by comparing them with the OpenLLM benchmark (vi) and further experiments with the HELM benchmark (vii).
i) Holmes results are reliable.
@@
Figure 8 shows the reliability of probing-based evaluations using averaged results across random seeds and LMs. Single outliers are datasets that are too hard for all LMs, either because the sample size is too small or the linguistic phenomena under test are too complex. First, a low average deviation ($\sigma=0.02$) across five seeds underscores the reliability of probing-based measures. These results also highlight the stability of probing results over prompting-based evaluations, where prompt paraphrasing leads to deviations of $\sigma=0.07$ reported in Mizrahi et al. (2024). Next, substantial compression (average $I=1.9$) and selectivity (average $S=0.31$) further confirm the probes’ reliability. Note, for selectivity, we consider only base-sized model (10m-200m parameters) for computational efficiency. Interestingly, two parallel trends emerge. More challenging datasets with many labels, like POS tagging, are arranged around a selectivity of 0.1 to 0.4 and a task metric of 0.3. In contrast, for easier binary classification tasks (such as linguistic applicability), we observe selectivity around 0.2 to 0.5 and a task metric of 0.6 to 0.9. Furthermore, our analysis reveals a statistically significant positive correlation ($p<0.05$) between the task metrics and both compression ($\tau=0.64$) and selectivity ( $\tau=0.65$). This finding provides strong evidence for the reliability of our task metric, thereby justifying its use as the primary evaluation measure in our study.
@@
Figure 8: Reliability evaluation Holmes results to ensure low deviation across random seeds, high information compression (log), and high selectivity. Every dot represents the averaged results of one probing dataset across LMs. The x-axis represents the task metrics (either person correlation or macro $F_{1}$).
ii) LMs’ linguistic competence is manifold.
@@
We focus on what Holmes tells us in general and regarding formal and functional phenomena, as defined in § 2. We report in Figure 9 the task metric, discriminability, and selectivity, averaged for every phenomena type. Note, discriminability (Rodriguez et al., 2021) quantifies the alignment of LMs ranking of one specific dataset compared to the overall rankings using the Kendall Tau correlation. Considering these three metrics, all tested LMs strongly encode formal phenomena (morphology and syntax), which often depend on the local neighborhood of words. Therefore, we assume that LMs approximate these co-occurrences during pre-training with high precision. For example, the specific POS tag of a word, like man (noun), primarily depends on its surroundings, such as the frequent predecessor the. In contrast, LMs encode less information about functional phenomena (semantics, reasoning, and discourse) since they show a relatively low performance regarding the task metric. For these functional phenomena, we assume more complex co-occurrences are required to capture the broad context in language, such as the rhetorical relation of two distant text spans. Despite these differences between formal and functional phenomena types, they contribute to the benchmark in a balanced way. A low to medium discriminability indicates that none of these linguistic phenomenon types dominates the overall LM rankings.
@@
This balanced influence of the five phenomena types is further visible when considering their ranking correlations (Figure 10, left). A high average correlation of $68.4\pm 7.5$ with the overall results (last column/row) hints that they are facets of a broader occurrence but share common characteristics. Still, breaking into categories is meaningful, as the phenomena types (first five columns/rows) are medium correlated (average of $54.7\pm 13.9$). Analyzing the results of phenomena types further highlights the value of this distinction. While results of semantics and reasoning are similarly correlated with the overall results ($73.9$ and $75.6$), their direct correlation ($58.4$) indicates their supplementary nature. Further, discourse results show the lowest correlation with others ($44.4\pm 14.7$), indicating a particular scope.
@@
iii) Encoder architecture equips LMs with high linguistic competence.
Next, we discuss the impact of model architecture on the linguistic competence of LMs. In Figure 11 (left), we compare encoder and decoder LMs. Due to the absence of big encoder LMs, we consider five encoder and six decoder LMs with up to 220m parameters. Encoder LMs show a higher $mwr$ of $52\%$ than decoder LMs ($21\%$). This observation is the most saturated for morphology or syntax, encompassing a variety of token-level phenomena, like part-of-speech. We assume that the missing bi-directional encoding of decoder LMs causes this lower performance because the available context of one token heavily depends on its position. Thus, even common tokens, like the, have different potential representations - at the beginning or middle of a sentence. These instabilities are further evident when considering Figure 11 (right) which reports the accuracy for the top-20 most common POS tokens (such as the) based on the pos, xpos, upos dataset. Given their high frequency, one expects stable prediction performance. Surprisingly, encoder LMs (BERT and RoBERTa) show higher median accuracy and lower deviations compared to the same-size decoder counterpart (GPT2). While scaling model size to 12B (Pythia) and 70B (Llama-2) allows for improved accuracy and lower deviations, decoder LMs do not match the encoder performance, even up to 700 times bigger.
Figure 10: Kendall-tau correlation within Holmes (left) and compared to OpenLLM (right). Green stars indicate significant correlations ($p<0.05$).
@@
iv) More parameters improve LMs’ linguistic competence.
We discuss how the number of parameters influences the linguistic competence of LMs. Given the variety of LMs of different sizes, we focus on the Pythia (decoder-only) and T5 (encoder-decoder) families. From Figure 12, we observe for both Pythia and T5 that the linguistic competence scales with model size, and it is particularly pronounced after exceeding 0.5B (Pythia) and 1.0B (T5) parameters. Again, model architecture is crucial, as T5 LMs (encoder-decoder) exhibit a clearly higher mean winning rate of $40-70\%$ than Pythia (decoder-only) ones with $mwr$ of $20-60\%$. Further, we found formal phenomena evolving differently with increased model size than functional ones. Specifically, morphology and syntax start at a lower level, with an apparent performance jump after 0.5B (Pythia) and 1.0B (T5) parameters, followed by slow but steady growth. Differently, semantics, reasoning, and discourse start at a higher $mwr$, followed by a continuous improvement as the model size grows. From these results, we assume that more parameters enable language models to better approximate simple word co-occurrences in nearby contexts. While handling formal phenomena like word dependencies, they struggle with more distant and complex co-occurrences, such as rhetorical relations.
@@
Tülu-2 +6% +3% -13% +1% -13% -4%
Orca-2 0% -4% -6% +3% -2% -3%
Llama-2-chat +9% +6% 0% +7% +1% +4%
Vicuna-v1.5 +26% +9% 0% +8% +2% +7%
@@
Tülu-2 +14% 0% -9% -4% +1% -2%
Llama-2-Chat +24% +13% +3% +3% +13% +10%
Average +10% +5% -3% +4% -1% +2%
@@
Table 2: The mixed effect of instruction tuning on the mean winning rate compared to the pre-trained LMs.
v) Instruction-tuned LMs get better at mimicking language than understanding it.
@@
We focus on how instruction tuning affects LMs’ linguistic competence and compare tuned and pre-trained LMs, for example, FLAN-UL2 vs. UL2. Table 2 shows less saturated effects for the overall scope while being more pronounced for the five phenomenon types - again emphasizing the structured and comprehensive evaluation of linguistic competence. On average, we found instruction tuning has the highest effect on morphology ($+10\%$) followed by syntax ($+5\%$), reasoning ($+4\%$), and a negative effect for semantics $-3\%$ and discourse $-1\%$. These results confirm previous assumptions that instruction tuning updates are often superficial (Yadav et al., 2023; Hershcovitch et al., 2024; Sharma et al., 2023) and that LMs get better at mimicking language (formal phenomena) than understanding it, measured with functional phenomena (Mahowald et al., 2024). Further, larger models benefit more from instruction tuning. Llama-2-70b-Chat and FLAN-UL2 gain up to $+24\%$ and $+41\%$ for morphology and $+10\%$ and $+12\%$ on average. When comparing LMs based on Llama-2-13B, we see that specific fine-tuning methods shape the LMs differently. The top-ranked 13B LM for Holmes and OpenLLM, Vicuna, was trained on fewer instructions than others (125k) but of higher quality. This high quality seems important as LMs with more instructions but lower quality (Tülu with approx. 330k instructions) lose performance, the same for 70B versions. Further and aligned with the previous comparison with OpenLLM results, reasoning specialization (Orca-2) is reflected in the corresponding phenomena. These insights show again that while providing a particular perspective, Holmes shows apparent differences between LMs and allows us to map them to methodological decisions.
vi) Internals of LMs are partly aligned with their linguistic performance.
We analyze the alignment of the probing-based LM rankings of Holmes with prompting-based ones when evaluating downstream using the LMs responses (linguistic performance). Specifically, we compare against OpenLLM (Beeching et al., 2023). [^4] Unlike other benchmarks like HELM (Liang et al., 2023), OpenLLM covers many open LMs leading to high overlap with Holmes. Figure 10 (right) shows Holmes and OpenLLM rankings of jointly evaluated LMs are medium correlated, hinting that LMs’ linguistic competence is partly reflected in their language utterances when solving concrete tasks. While syntax, semantics, and discourse show similar correlation (54.7 to 58.0), morphology and reasoning exhibit a substantially higher one of 65.3 and 77.5. These results suggest that LMs’ reasoning abilities are reflected in their internal representations when evaluating related phenomena like identifying the cause of negations. These correlation patterns are consistent across the three most meaningful OpenLLM datasets (MMLU, TruthfulQA, and GSM8K). As TruthfulQA shows lower correlations with the linguistic phenomena and other datasets within OpenLLM, we presume this dataset captures distinctly different skills (possibly knowledge).
vii) Prompting is not a substitute for probing when evaluating LMs’ linguistic competence.
Finally, we compare probing- and prompting-based LM rankings on the jointly evaluated BLiMP tasks (Warstadt et al., 2020) of Holmes and HELM (Liang et al., 2023). Results (Appendix, Figure 15) show apparent discrepancies (rank correlation $\tau=0.05$) when evaluating LMs’ internal representations or their responses (linguistic performance) to HELM instructions. As most prompting-based results from HELM fall below the random baseline, only probing-based evaluation can effectively isolate the assessment of linguistic phenomena. In contrast, prompting-based methods mix this assessment with other abilities, such as instruction following. Similar to Hu and Levy (2023), these insights show the need for a more comprehensive comparison of different evaluation protocols like probing, prompting, or log-probabilities (used in HELM in Figure 33 on page 58 as a workaround for BLiMP). Nevertheless, probing provides a unified evaluation protocol assessing the diversity of linguistic phenomena using representations of tokens, spans, or whole texts beyond minimal pair tasks testing whether correct or wrong sentences are preferred.